OpenStudy (anonymous):

is 3 a irrational number

sammixboo (sammixboo):

Is it a never ending decimal?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so the answer is what?/

sammixboo (sammixboo):

Is 3 a never ending decimal or not?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 is a whole number

OpenStudy (anonymous):

is 1 over 8 an irrational number

sammixboo (sammixboo):

No.. 3 isn't even a never ending decimal. Like I replied in one of your other questions. An irrational number is a real number that cannot be written as a simple fraction aka a never ending fraction

sammixboo (sammixboo):

Tell me what is \(\sf \large \frac{1}{8}\) is?

sammixboo (sammixboo):

(You take 1 and divide it by 8)

OpenStudy (anonymous):

0.125

sammixboo (sammixboo):

Can 0.125 be written as a fraction?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is 1.24562 a rational number or irrational?

sammixboo (sammixboo):

Stop hold on a second please

OpenStudy (anonymous):

you get it

sammixboo (sammixboo):

0.125 can be written as a fraction... All you got to do is put 125 over 1,000 \(\sf \large ->\) \(\sf \large \frac{125}{1,000}\) 0.125 = \(\sf \large \frac{125}{1,000}\) What is \(\sf \large \frac{125}{1,000}\) simplified?

OpenStudy (anonymous):

0.125

sammixboo (sammixboo):

How did I get \(\sf \large \frac{125}{1,000}\)? All I did was take the decimal, 0.125, and put it over 1, \(\sf \large \frac{0.125}{1}\), and then multiply both the numerator and denominator by three 10's.. (numerator) \(\sf \large 0.125*10*10*10~=~125\) (denominator) \(\sf \large 1*10*10*10~=~1,000\) Vuala, and I meant what is \(\sf \large \frac{125}{1,000}\) simplified as a fraction?

OpenStudy (anonymous):

can you just give me the answer please!!

sammixboo (sammixboo):

All you gotta do is divide 125 by 125 and 1,000 by 125 \(\sf \large \frac{125/125}{1,000/125}\)

sammixboo (sammixboo):

And this will help you with future problems

OpenStudy (anonymous):

0.008

OpenStudy (anonymous):

and just 8